Solve the system of equations by graphing on a sheet of paper, and then state the solution. Explain how you solved. Thank You!!!!!
y=2x-1
6x-y=13

Answer by plourde216(7)

The first thing that you need to do is to solve the second equation for y. This will put the equation into slope-intercept form.
6x-y=13 subtract 6x from both sides,
-y=13-6x divide both sides by -1 (to make the solution y= not -y=)
y=-13+6x
Now you have 2 equations in slope intercept form. You need to graph the 2 equations by hand or with the aid of a graphing calculator. The solution is going to be where the 2 graphs intersect. In this case, (3,5). Whenever you get your answer, be sure to substitute this ordered pair back into your original equations. This allows you to check your work.
6x-y=13 y = 2x-1
6(3)-5=13 5 = 2(3-1)
13=13 This one checks out 5=5
Because the solution checks out in both equations, you have found the correct solution.